how much would you need to deposit in an account now in order to have $6000 in
the account in 15 years? Assume the account earns 8% simple interest. Round
your answer to the nearest cent

Answer by stanbon(75887)

how much would you need to deposit in an account now in order to have $6000 in
the account in 15 years? Assume the account earns 8% simple interest. Round
your answer to the nearest cent
---------------------
A(t) = P(1+r*t)
------
6000 = P(1+0.08*15)
------
6000 = P(1+1.20)
-----
P = 6000/(2.2)
Principal = $2727.27
--------------
Cheers,
Stan H.
------------